OpenStudy (anonymous):

calculate the volume of a cone with height 5 ft and diameter 4 ft

OpenStudy (anonymous):

Volume = \[\frac{ 1 }{ 3 }*\pi*r^2*h\] where r = radius of cone and h = height of the cone remember diameter = 2*radius

OpenStudy (anonymous):

Thank you